### v2

@@ 2 Methods
2.1 Ensemble LLMs for Topic Labeling
In this section, we present the detailed methodologies for the four stages of the ensemble framework illustrated in Figure 1: (1) topic labeling by each selected open-source LLM; (2) relevancy scoring based on the embedding distance between each LLM’s reasoning and the topic description; (3) agreement evaluation among LLMs and the removal of potential outliers; and (4) the ensemble of scores and labels from the final set of LLMs.
Figure 1: The ensemble framework is illustrated in two parts. The upper panel depicts the pipeline for topic labeling and relevancy scoring. Each LLM is prompted with text and a detailed topic description, and returns both a dichotomous label and related phrases identified from the text. Relevancy scores are calculated using the cosine similarity between the embeddings of the topic description and the returned phrases. The lower panel represents the ensemble methodology, which leverages the heterogeneous performances of diverse open-source, deployable-sized language models. The final ensemble of labels is determined by removing false positives from the union of positive cases, using the optimal threshold on the ensemble of relevancy scores that predicts the intersection of positive cases with the highest F1-score.
@@ 2.1.1 Topic Labeling
A collection of open-source language models should be used to conduct the same labeling and scoring task for a topic of interest. In this experimentation, we selected four small-sized (7-8 billion parameters) language models that are open-source on huggingface: Llama-3.1-8B-Instruct [34], Qwen2.5-7B-Instruct [50], Mistral-7B-Instruct-v0.3 [51], and Vicuna-7b-v1.5 [48, 52]. Each model can be deployed in a secure computing environment and is recognized for its capacity in text generation inference [53]. To determine the presence of a topic in a given textual data, we prompted each LLM with the text and a detailed description of the topic. The prompt design is shown in the upper panel of Figure 1. Although techniques such as few-shot prompting [54], chain-of-thought [55], and other prompt engineering techniques [56, 57] can be employed, this study focuses on comparing the ensemble performance with individual performances of LLMs in topic labeling tasks. We used the same prompt scheme across all LLMs, allowing diverse models to reveal their heterogeneous capacities in performing the same task.
@@ 2.1.2 Relevancy Scoring
Recent transformer-based embeddings [58], exemplified by BERT (Bidirectional Encoder Representations from Transformers) models [59], mark a significant advancement in capturing contextualized word representations by simultaneously enlabeling words and their relationships within a sentence as numerical embedding vectors [60]. The distance between two sentences can be quantified using the cosine similarity between their vectors in the embedding space [61].
In our prompt design, each LLM is asked to return both a dichotomous label indicating the presence of a given topic in the text, as well as the topic-related phrases identified within the text, which reflect the LLM’s reasoning. Subsequently, a relevancy score is calculated using the cosine similarity between the embeddings of the topic description and the related phrases returned by an LLM. This score measures the extent to which the evidence provided by an LLM is relevant to the queried topic and reflects how explicitly the text discusses the topic.
In our experiments, embeddings are generated using the pre-trained Sentence-BERT model all-mpnet-base-v2 [60], which maps sentences (truncated to a maximum of 384 words) into a 768-dimensional vector space. This model, trained on over 1 billion sentences, has demonstrated optimal performance in semantic search according to the massive text embedding benchmark leaderboard [62].
Lastly, for a given topic description, we calculated a baseline cosine similarity score between the description and an empty string. We set the observed scores under the baseline to zero, and subtracted the baseline score from the observed score, yielding a final relevancy score ranging from 0 to 1.
@@ 2.1.3 Agreement Evaluation
Before ensembling, we conducted an agreement evaluation on the labels and scores returned by the LLMs. Gwet’s AC1 [63] and Fleiss’s Kappa [64] are the preferred inter-rater reliability metrics in scenarios where the topic of interest is rarely prevalent [16, 64]. Other metrics, such as Cohen’s Kappa [65], may result in the “high agreement, low reliability” paradox in rare event labeling [66, 64, 67]. To evaluate the agreement among the continuous relevancy scores across LLMs, we categorized the scores into ten ordinal levels, each spanning 0.1, before calculating unweighted Gwet’s AC1 and Fleiss’s Kappa.
Although the ensemble of LLMs should benefit from the heterogeneity of diverse LLMs, those whose labels and scores significantly deviate from the rest may detrimentally affect the ensembling process. Thus, to remove potential outlier LLMs, we calculated the increase in Gwet’s AC1 when each LLM was excluded from the full collection. If this increase surpasses a predetermined threshold, the LLM should be excluded from the ensembling candidates. The choice of threshold depends on the needs of specific applications, and in our empirical evaluation, we required the increase to surpass 10% of the overall AC1.
@@ 2.1.4 Ensembling
From each LLM, we retrieve a binary label and a relevancy score ranging from 0 to 1 if the label is positive. An ensemble of relevancy scores can be generated using the principal component that explains the most variability in the data, through principal component analysis (PCA), provided that the sample size of the text data is sufficient. Then, the PCA ensemble of scores is used to inform the ensembling of labels.
The primary objective in ensembling the labels is to eliminate potential false positives from the union of positive labels returned by the LLMs. The union of positive labels from multiple LLMs inherits all the false positives from each LLM. Conversely, the intersection of positive labels, defined as those agreed upon by more than half of the LLMs, could be too conservative, potentially failing to include all true positives. Thus, the ensembling of labels involves finding a balance between the union and the intersection, guided by the PCA ensemble of relevancy scores. Specifically, we examined each threshold on the scale of the PCA ensemble of scores and identified the optimal threshold that predicts the intersection of positive cases with the highest F1-score. This threshold represents the optimized trade-off between sensitivity and precision, particularly for topics with rare prevalence [68]. Subsequently, we excluded those cases in the union of positive cases whose scores fell below this threshold as false positives.
@@ 2.2 Empirical evaluation
In this study, we evaluated the ensembling framework through two case studies of psychological textual data, focusing on eating disorders. We used human annotation to evaluate the labeling performances of the four open-source LLMs previously listed, as well as the ensembles of all possible combinations of them. For the public dataset of Reddit posts, we also used closed-source large-sized GPT-4o as a judge [48]. The Reddit posts, along with a human-annotated subset, are made public through the University of Virginia Dataverse [V3/NENELT_2025].
@@ 2.2.1 Case Study 1: Overlapping Topics Between ED and diet culture on Reddit
Social media has shown to promote diet culture, characterized by health myths about food and eating, as well as a moral hierarchy of bodies [71, 72]. Though diet culture lacks a uniform definition [71], it is defined in this study as a system of beliefs that equates thinness with health and oppresses people who do not align with standards of body shape, size, and weight [72]. Research shows that dieting and other weight-control methods significantly increase the risk of harmful eating practices associated with disordered eating or ED [73]. In this case study, we aimed to identify potentially overlapping topics between ED and general dieting online discussions to better understand the similarities between diet culture and ED. Extended from previous analyses of ED online discussions [74, 75], we identified 15 candidate topics that are likely prevalent in both dieting and ED communities, as listed in Table 2.2.1 (a).
{tableorg}
Descriptions of topic of interests in two empirical case studies 0pt Topic Short Name Description Binge eating binge Binge eating. Body dissatisfaction bodyhate Body dissatisfaction, feel bad about body image and appearance. Calorie count calorie Count calorie. Food cravings crave Craving for high calorie food or carbs. Depressed mood depressedmood Depressed mood, feeling depressed. ED recovery ed Eating disorders(ED) diagnosis or recovery, ED includes anorexia nervosa, anorexic, bulimia, bulimic, binge eating disorders, arfid, osfed, pica. Physical exercise exercise Physical exercise. Fear foods fearfood Fear certain foods. Fear of weight gain feargain Fear of body weight gain. Must related to body weight. Weight gain gain Body weight gain. Weight loss loss Body weight loss. High protein diet protein High protein diet, carbohydrate-reduced(low-carb) high-protein diet. Relationships relation Family and social relationships. Restriction restrict Restrict nutrition or calorie intake. Ideal body image idealbody Ideal body image including thinness, skinny body, low body fat and lean body mass. Case study 1: labeling overlapping topics between dieting and ED discussions among reddit posts 0pt Subcategories of Weight Stigma in Treatment Assumptions of health based on appearance or weight. Healthcare decisions based on weight. Treatment decision based on weight. Not sick enough, or not thin enought, or not low weight enough. Not taken seriously because of weight, Inadequate care based on weight. Diet promotion. Encouragement of weight loss. Reassurance of thinness. Reassurance that will not get fat in treatment. Negative attitudes, discrimination, or prejudice based on body weight or size. Weight blamed for health issues or concerns. Weight tied to personality characteristics. Case study 2: labeling weight stigma experienced by ed patients during treatment
We collected up to 1000 top-rated posts and up to 1000 newest posts per subreddit forum across three groups: dietary, fitness, and eating disorders, using Python package PRAW (7.7.1). Posts in dietary and fitness forums contain content of general dieting online discussions. In total, we collected 77,175 posts from 53,784 authors across 115 subreddit forums. Data descriptions and the designated searching keywords are shown in Table 1. For ED forums, we included 12 subreddits: ‘EatingDisorders’, ‘bulimia’, ‘intuitiveeating’, ‘AnorexiaNervosa’, ‘EDanonymemes’, ‘EDAnonymous’, ‘BingeEatingDisorder’, ‘fuckeatingdisorders’, ‘ARFID’, ‘eating_disorders’, ‘save_food’, ‘edsupport’. The data acquisition took place in February 2024.
As a public dataset, we conducted the evaluation in this case study using both human annotators and the large-size, closed-source GPT-4o. Firstly, for practicality, we randomly sampled 1,080 posts—10 posts per forum from 108 forums each containing at least 10 posts. One author labeled all 1,080 posts, and another author independently labeled a subset of 200 posts. The Gwet’s AC1 and Fleiss’ Kappa scores on the 200 posts labeled by the two coders were 0.972 (95% CI: 0.966 - 0.979) and 0.851 (95% CI: 0.817 - 0.886), respectively, indicating excellent agreement and reliability of the human annotation. Secondly, we used GPT-4o to label and rate all 77,175 posts. The Gwet’s AC1 and Fleiss’ Kappa with human annotation were 0.976 (95% CI: 0.974 - 0.979) and 0.885 (95% CI: 0.873 - 0.897), respectively, suggesting that large-sized LLMs can achieve close-to-human performance in these tasks, as previously recognized [76, 77].
@@
2.2.2 Case Study 2: ED Patients’ Experiences with Weight Stigma during Treatment
In the second case study, we used a protected clinical dataset of ED patients narratives about their experience with weight stigma during treatment. A total of 1,368 texts from patients treated for eating disorders were collected within a clinical setting. Two domain experts independently coded the dataset, determining the presence of weight stigma in patient narratives by identifying any of the subcategories listed in Table 2.2.1 (b). The Gwet’s AC1 and Fleiss’ Kappa between two coders’ annotation were 0.625 (95% CI: 0.583 - 0.667) and 0.541 (95% CI: 0.494 - 0.588), respectively. Due to the relatively low human inter-rater reliability, we used the subset of 1,080 texts where both coders agreed as the final human annotation to evaluate the labeling performance of LLMs and their ensembles.

### v4

@@ 2 Methods
2.1 Ensemble Framework
As shown in Figure 1, the ensemble framework of multiple LLMs is analogous to employing multiple human raters for rating curated topics in free-text data. In this section, we detail the key steps of the ensemble framework: (1) topic labeling by each selected open-source LLM; (2) relevance scoring based on the embedding distance between each LLM’s reasoning and the input topic description; (3) agreement evaluation across LLMs with the removal of potential outliers; and (4) ensembling of scores and labels from multiple LLMs.
@@ 2.1.1 Topic Labeling
A collection of open-source language models should be used to conduct the same labeling and scoring task for a topic of interest. In this experimentation, we selected four small-sized (7-8 billion parameters) language models that are open-source on huggingface: Llama-3.1-8B-Instruct [50], Qwen2.5-7B-Instruct [67], Mistral-7B-Instruct-v0.3 [68], and Vicuna-7b-v1.5 [65, 69]. Each model can be deployed in a secure computing environment and is recognized for its capacity in text generation inference [70]. The largest selected model requires a minimum of 20 GB full-precision VRAM, and our experiments were conducted on a single NVIDIA V100 GPU.
To determine the presence of target topics in a given text, we prompted each LLM with the text and detailed descriptions of the topics, using the following template:
Does the paragraph mention any of the following topics:
(1) {topic short name 1}: {description}.
(2) {topic short name 2}: {description}. ...
Return answer in format:
(1) {topic short name 1}: [yes/no], related phrases if any:
(2) {topic short name 2}: [yes/no], related phrases if any: ...
Paragraph: ‘{text}‘
Figure 2 illustrates an example of prompting a single LLM to determine whether the topics “body dissatisfaction (bodyhate)” and “afraid of body weight gain (feargain)” are present in a paragraph-length post from an eating disorder–related Reddit forum. The response from Llama-3.1-8B-Instruct consists of a positive label (“yes”) along with related phrases for each topic—e.g., “wear baggy clothes to hide what I look like” and “too scared to talk to anyone” were identified as relevant to body image dissatisfaction. Although techniques such as few-shot prompting [71], chain-of-thought [72], and other prompt engineering techniques [73, 74] can be employed, this study focuses on comparing the ensemble performance with individual performances of LLMs in topic labeling tasks. We applied the same prompt template across all LLMs, allowing the models to reveal their heterogeneity in performing the same task.
Figure 2: The procedure of topic labeling and relevance scoring performed by a single LLM, applied to an example text post from an eating disorders Reddit forum. Each LLM is prompted with the text, the target topics, and the corresponding detailed description of each topic. It then returns a dichotomous label indicating the presence of each topic, along with the text phrases identified as related to that topic, reflecting its reasoning. Then, a relevance score is computed for each topic by calculating the cosine similarity between the embedding of the topic description and the embedding of the returned phrases, thereby measuring the semantic relevance of that topic within the text.
@@ 2.1.2 Relevancy Scoring
Recent representation-learning–focused language models, exemplified by BERT (Bidirectional Encoder Representations from Transformers) and its variants [75, 36, 76], mark a significant advancement in capturing contextualized word, phrase, and sentence representations. These models generate numerical embedding vectors that encode the meaning of words in a given text, their contextual relationships, and the overall semantic context. The semantic distance between two texts (sentences or phrases) can be quantified using the cosine similarity between their vectors in the embedding space [77].
Since each LLM is prompted to return topic-related phrases identified within the input text, we calculate a relevancy score for each topic by computing the cosine similarity between the embedding of the topic description and the embedding of the phrases returned by the LLM. In the example shown in Figure 2, for the topic body dissatisfaction, the topic description provided in the prompt is “body dissatisfaction, feel bad about body image and appearance.” The cosine similarities between this description and the LLM-returned phrases are 0.43 for “wear baggy clothes to hide what I look like” and 0.22 for “too scared to talk to anyone.” By taking the maximum, the final relevancy score of topic bodyhate in this post is 0.43 for this LLM. This score measures the extent to which the evidence provided by an LLM is relevant to the queried topic and reflects how explicitly the topic is discussed in the input text.
Note that cosine similarity between two texts can differ across embedding models potentially due to the anisotropy of the embedding space [78, 79]; some models produce similarities averaging around 0.5, while some yield averages closer to 0. For the generalizability of our framework across different embedding models, we calculate a baseline cosine similarity score between each topic description and an empty string, aiming to recalibrate similarity so that the distance to an empty string is centered at zero. Specifically, we adjust the similarity scores between the topic description and the returned phrases by subtracting this baseline and clamping negative values to zero, yielding a final relevancy score in the range [0, 1].
In our experiments, embeddings are generated using the pre-trained Sentence-BERT model all-mpnet-base-v2 [76], which maps long text or sentences (truncated to a maximum of 384 words) into a 768-dimensional vector space. This model, trained on over 1 billion sentences, has demonstrated optimal performance in semantic search according to the massive text embedding benchmark leaderboard [80].
@@ 2.1.3 Agreement Evaluation
Before ensembling, we conducted an agreement evaluation on the labels and scores returned by the LLMs. Gwet’s AC1 [81] is a preferred inter-rater reliability metric in scenarios where the topic of interest is rarely prevalent [15]. Other metrics, such as Cohen’s Kappa [82] and Fleiss’s Kappa [83], can result in the “high agreement, low reliability” paradox in rare-event labeling when the topic of interest is present at low prevalence [84, 83, 85]. To evaluate the agreement among the continuous relevancy scores across LLMs, we categorized the scores into ten ordinal levels, each spanning 0.1, before calculating unweighted Gwet’s AC1. We also reported Fleiss’s Kappa, as it is a commonly used inter-rater reliability coefficient for more than two raters (or LLMs in our case). The two coefficients are expected to differ because Gwet’s AC1 employs an adjusted chance agreement calculation to account for imbalances in category prevalence. Further details on Gwet’s AC1 and Fleiss’s Kappa are provided in Appendix A.
Although the ensemble of LLMs should benefit from the heterogeneity of diverse LLMs, those whose labels and scores significantly deviate from the rest may detrimentally affect the ensembling process. Thus, to remove potential outlier LLMs, we calculated the increase in Gwet’s AC1, which is less sensitive to low topic prevalence, when each LLM was excluded from the full collection. If this increase exceeds a predetermined threshold (in our experiments, greater than 10% of the overall AC1), the LLM is excluded from the ensembling candidates.
@@ 2.1.4 Ensembling
From each LLM, we retrieve a binary label and a relevancy score ranging from 0 to 1 if the label is positive. As shown in Figure 1, we propose a simple ensembling procedure that consists of two stages: (1) ensemble of relevancy scores, and (2) ensemble of dichotomous labels.
Figure 3: Process of the ensemble of labels by multiple LLMs. Relevancy scores from multiple LLMs are aggregated using PCA to form a single ensemble score. By sweeping thresholds on this scale, sensitivity, precision, and F1-score are computed against the intersection label (which is determined by the majority voting). The optimal threshold maximizes F1 and positive texts with relevancy under the threshold are considered potential false positives and are reassigned as negatives in the final label.
The ensemble of the relevancy scores aims to calculate a single relevancy of a topic in a text, aggregated across non-outlier LLMs. Compare to an unweighted average of scores across multiple LLMs, principal component analysis (PCA) looks for the linear combination of variables that explains the most variance in the data. For example, the principal component of four score variables (vectors of $N$ texts) $x_{i}$ from four LLMs is given by $\text{PC}_{1}=w_{1}x_{1}+w_{2}x_{2}+w_{3}x_{3}+w_{4}x_{4},$ where the weights $w_{i}$ are chosen to maximize the variance of $\text{PC}_{1}$ subject to the normalization constraint $\sum_{i=1}^{4}w_{i}^{2}=1$. By achieving maximum variability in $\text{PC}_{1}$, we calibrate texts with mostly higher relevancy scores from LLMs to the upper end of the single ensemble score scale, and texts with mostly lower scores to the lower end.
For the ensemble of labels, we detailed the process in Figure 3. The primary objective in ensembling the labels is to eliminate potential false positives from the union of positive labels returned by the LLMs. The union of positive labels from multiple LLMs inherits all the false positives from each LLM. Conversely, the intersection of positive labels—defined as those agreed upon by more than half of the LLMs, following the majority-voting convention commonly used in human rating processes—may serve as a closer proxy to the true label, but can potentially fail to include all true positives. We term the union label as the resulting label after taking the union of positive labels from multiple LLMs, and the intersection label as the resulting label after taking the intersection (majority voting) of positives from multiple LLMs. The ensembling of labels involves finding a balance between the union and the intersection (by majority voting), guided by the PCA ensemble of relevancy scores. Specifically, we examined each threshold on the scale of the PCA ensemble of scores and identified the optimal threshold that predicts intersection labels with the highest F1-score. For example, in Figure 3, at each threshold on the scale of the PCA ensemble score, we compute a predicted label by classifying texts below the threshold as negative and those above as positive. We then calculate sensitivity and precision between the predicted label and the intersection label, defined as
$\text{Sensitivity}=\frac{\text{TP}}{\text{TP}+\text{FN}},\quad\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}},$
where TP, FP, and FN denote true positives, false positives, and false negatives, respectively. As the threshold moves from low to high on the PCA ensemble score scale, recall decreases while precision increases. We identify an optimal threshold where the F1-score (the harmonic mean of precision and recall),
$F1=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}},$
is maximized. This threshold represents the optimized trade-off between sensitivity and precision, particularly for topics with rare prevalence [86]. Lastly, we modify the union label into a final ensemble label as follows: texts in the union of positive cases whose scores fall below the optimal threshold are considered potential false positives and are reassigned as negatives in the final label.
@@ 2.2 Empirical Evaluation
In this study, we evaluated the ensembling framework through two case studies of psychological textual data focused on eating disorders: (1) a public dataset of Reddit posts discussing topics related to dieting and ED, and (2) a private (protected) dataset describing ED patients’ experiences of weight stigma during treatment. We used human annotation to evaluate the labeling performances of the four open-source LLMs previously listed, as well as the ensembles of all possible combinations of them. For the public dataset of Reddit posts, we also used closed-source large-sized GPT-4o as a judge [65] to evaluate their labeling performances. The Reddit posts, along with a human-annotated subset, are made public through the University of Virginia Dataverse [89]. Note that human annotation is only used in evaluating the labels produced by individual LLMs or their ensembles; the ensemble framework does not rely on human evaluation.
The study of weight stigma was approved by the Institutional Review Board (IRB) of University of Louisville (Protocol 20.0837). The study of Reddit posts used publicly available, de-identified data and was approved by the Institutional Review Board for the Social and Behavioral Sciences (IRB-SBS) of University of Virginia (Protocol 7572).
Table 1: Descriptions of topic of interests in two empirical case studies
Topic Short Name Description
Binge eating binge Binge eating.
Body dissatisfaction bodyhate Body dissatisfaction, feel bad about body image and appearance.
Calorie count calorie Count calorie.
Food cravings crave Craving for high calorie food or carbs.
Depressed mood depressedmood Depressed mood, feeling depressed.
ED recovery ed Eating disorders(ED) diagnosis or recovery, ED includes anorexia nervosa, anorexic, bulimia, bulimic, binge eating disorders, arfid, osfed, pica.
Physical exercise exercise Physical exercise.
Fear foods fearfood Fear certain foods.
Fear of weight gain feargain Fear of body weight gain. Must related to body weight.
Weight gain gain Body weight gain.
Weight loss loss Body weight loss.
High protein diet protein High protein diet, carbohydrate-reduced(low-carb) high-protein diet.
Relationships relation Family and social relationships.
Restriction restrict Restrict nutrition or calorie intake.
Ideal body image idealbody Ideal body image including thinness, skinny body, low body fat and lean body mass.
(a) Case study 1: labeling overlapping topics between dieting and ED discussions among reddit posts
Subcategories of Weight Stigma in Treatment
Assumptions of health based on appearance or weight.
Healthcare decisions based on weight. Treatment decision based on weight.
Not sick enough, or not thin enough, or not low weight enough.
Not taken seriously because of weight, Inadequate care based on weight.
Diet promotion.
Encouragement of weight loss.
Reassurance of thinness. Reassurance that will not get fat in treatment.
Negative attitudes, discrimination, or prejudice based on body weight or size.
Weight blamed for health issues or concerns.
Weight tied to personality characteristics.
(b) Case study 2: labeling weight stigma experienced by ed patients during treatment
@@ 2.2.1 Case Study 1: Overlapping Topics Between ED and Diet Culture on Reddit
Social media has shown to promote diet culture, characterized by health myths about food and eating, as well as a moral hierarchy of bodies [90, 91]. Though diet culture lacks a uniform definition [90], it is defined in this study as a system of beliefs that equates thinness with health and oppresses people who do not align with standards of body shape, size, and weight [91]. Research shows that dieting and other weight-control methods significantly increase the risk of harmful eating practices associated with disordered eating or ED [92]. In this case study, we aimed to identify potentially overlapping topics between ED and general dieting online discussions to better understand the similarities between diet culture and ED. Extended from previous analyses of ED online discussions [93, 94], we identified 15 candidate topics that are likely prevalent in both dieting and ED communities, as listed in Table 1 (a).
@@
We collected textual posts from 115 Reddit forums across three topical groups—dietary, fitness, and eating disorders—using the Python package PRAW (v7.7.1). For each forum, we retrieved 1000 top-rated posts and 1000 newest posts, retaining unique entries. Posts in dietary and fitness forums contain content of general dieting online discussions. In total, we collected 77,175 posts from 53,784 authors across 115 subreddit forums. Data descriptions and the designated searching keywords are provided in Table 2. For ED forums, we included 12 subreddits: ‘EatingDisorders’, ‘bulimia’, ‘intuitiveeating’, ‘AnorexiaNervosa’, ‘EDanonymemes’, ‘EDAnonymous’, ‘BingeEatingDisorder’, ‘fuckeatingdisorders’, ‘ARFID’, ‘eating_disorders’, ‘save_food’, ‘edsupport’. The data acquisition took place in February 2024.
As a public dataset, we conducted the evaluation in this case study using both human annotators and the large-size, closed-source GPT-4o. Firstly, for practicality, we randomly sampled 1,080 posts—10 posts per forum from 108 forums that contains at least 10 posts. Two authors independently labeled all topics across the 1,080 posts. The Gwet’s AC1 and Fleiss’ Kappa scores on the 1,080 posts labeled by the two coders were 0.972 (95% CI: 0.966 - 0.979) and 0.851 (95% CI: 0.817 - 0.886), respectively, indicating excellent agreement and reliability of the human annotation. Secondly, we used GPT-4o to label and rate all 77,175 posts. The Gwet’s AC1 and Fleiss’ Kappa with human annotation were 0.976 (95% CI: 0.974 - 0.979) and 0.885 (95% CI: 0.873 - 0.897), respectively, suggesting that large-sized LLMs can achieve close-to-human performance in these tasks, as previously recognized [95, 96].
2.2.2 Case Study 2: ED Patients’ Experiences with Weight Stigma
In the second case study, we used a protected clinical dataset of ED patients narratives about their experience with weight stigma during treatment. A total of 1,368 texts from patients treated for eating disorders were collected within a clinical setting. Two domain experts independently coded the dataset, determining the presence of weight stigma in patient narratives by identifying any of the subcategories listed in Table 1 (b). The Gwet’s AC1 and Fleiss’ Kappa between two coders’ annotation were 0.625 (95% CI: 0.583 - 0.667) and 0.541 (95% CI: 0.494 - 0.588), respectively. Due to the relatively low human inter-rater reliability, we used the subset of 1,080 texts where both coders agreed as the final human annotation to evaluate the labeling performance of LLMs and their ensembles.